# Counterfactual Inference under Thompson Sampling

### v1

@@ 1. Introduction & Motivation
When framing the recommendation task as a decision-making problem (Jeunen et al., 2022; Jeunen and Goethals, 2023), it is natural to consider applications of causal or counterfactual inference (Wang et al., 2020; Gao et al., 2024; Bottou et al., 2013; Saito and Joachims, 2021). Methods from these adjacent fields can be used to answer general aggregate-level questions of the form: “What would the distribution of $R$ have been, had we deployed this alternative decision-making policy instead?” (Vasile et al., 2020)
The most common embodiment of such questions arises when carrying out offline evaluation tasks, aimed at estimating (differences in (Jeunen and Ustimenko, 2024a)) reward metrics for potential alternative recommendation models (Yang et al., 2018; Jeunen, 2019; Gilotte et al., 2018; Gruson et al., 2019; Jeunen et al., 2024). Additional to performance evaluation, these methods open up avenues for learning and fairness in recommendation—see Joachims et al. (2021) for a broad overview.
@@
Our work aims to close this gap. We derive a general expression for action propensities under TS, reducing to an integral that involves the probability density functions (PDFs) and cumulative distribution functions (CDFs) for the action-specific reward distributions. For the most common settings of normal distributions for continuous outcomes and Beta distributions for binary outcomes, we show that this integral reduces to an efficiently computable multivariate normal CDF or an analytically computable expression, respectively. We show that, under a common linearity assumption on the last layer of the reward model (Chapelle and Li, 2011; Watson et al., 2021), a normal distribution on model parameters directly translates to a distribution over model outcomes (i.e. rewards) (Zhang et al., 2021), enabling efficient propensity calculations. These novel derivations open up a range of potential use-cases where either or both the logging and target policies make use of TS or related action-selection procedures. We empirically validate our theoretical contributions by showcasing the efficacy of offline evaluation methods that would not have been feasible before, through reproducible experiments leveraging the Open Bandit Pipeline (OBP) (Saito et al., 2021).
@@ 2.1. Counterfactual or Off-Policy Estimation
Whilst this estimator is unbiased, its variance is often problematic, and a large literature on off-policy estimation has focused on variance reduction. Some trade in variance for a small bias (Ionides, 2008; Swaminathan and Joachims, 2015), others leverage additive control variates (Gupta et al., 2024b; Dudík et al., 2014; Vlassis et al., 2019; Farajtabar et al., 2018; Su et al., 2020; Jeunen and Ustimenko, 2024a).
Several related works have leveraged variants of these estimators to achieve empirical success in recommender systems evaluation (Yang et al., 2018; Gilotte et al., 2018; Gruson et al., 2019; Jeunen et al., 2024), as well as learning (Jeunen and Ustimenko, 2024b; Jeunen and Goethals, 2021a; Jeunen et al., 2020; Chen et al., 2019; Ma et al., 2020).
@@ 3. Methodology & Contributions
Intuitively, for every possible value of $r_{1}$, the reward estimates for all other actions need to be lower for $a_{1}$ to be selected. In the general case, this integral does not have a simple analytical solution. In what follows, we derive efficiently computable or closed-form solutions for common families of reward distributions: normal distributions for continuous outcomes, and Beta distributions for binary rewards.
@@ 3.1. Continuous Rewards: Normal Distributions
(6) and $\text{and~{}}\bm{V}=\begin{bmatrix}\sigma_{2}^{2}+\sigma_{1}^{2}&\dots&\sigma_{1}^{2}\\ \vdots&\ddots&\vdots\\ \sigma_{1}^{2}&\dots&\sigma_{n}^{2}+\sigma_{1}^{2}\\ \end{bmatrix}$
Note that this multivariate Gaussian CDF is an integral without a known analytical expression in and of itself—but one that is widespread, and where many accurate and efficient implementations are available to evaluate it (Virtanen et al., 2020).
@@ 3.1.1. Uncertainty on Model Parameters
Traditional TS places uncertainty on model parameters instead of model output. That is, instead of directly sampling outputs as $\hat{r}_{a}\sim f_{\theta}(a,x)$u we sample model parameters $\tilde{\theta}\sim\mathsf{P}(\Theta)$ and simply compute reward estimates as $\hat{r}_{a}=f_{\tilde{\theta}}(a,x)$, before picking the maximal one. It should be noted that in many common practical instantiations, these are equivalent. If we assume the parameters over which to place uncertainty $\theta\in\mathbb{R}^{d}$ to correspond to weights in a linear model and $f(a,x)$ is a fixed model transformation (e.g. a neural network, or the identity function in the case of a linear model), we have:
@@ Acknowledgements.
The author would like to thank Duane Rich for valuable comments on earlier versions of this manuscript.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the Nineteenth ACM Conference on Recommender Systems; September 22–26, 2025; Prague, Czech Republicbooktitle: Proceedings of the Nineteenth ACM Conference on Recommender Systems (RecSys ’25), September 22–26, 2025, Prague, Czech Republicdoi: 10.1145/3705328.3748011isbn: 979-8-4007-1364-4/2025/09
@@ 1. Introduction & Motivation
When framing the recommendation task as a decision-making problem (Jeunen et al., 2022; Jeunen and Goethals, 2023), it is natural to consider applications of causal or counterfactual inference (Wang et al., 2020; Gao et al., 2024; Bottou et al., 2013; Saito and Joachims, 2021; Jeunen, 2021). Methods from these adjacent fields can be used to answer general aggregate-level questions of the form: “What would the distribution of $R$ have been, had we deployed this alternative decision-making policy instead?” (Vasile et al., 2020)
The most common embodiment of such questions arises when carrying out offline evaluation tasks, aimed at estimating (differences in (Jeunen and Ustimenko, 2024)) reward metrics for potential alternative recommendation models (Yang et al., 2018; Jeunen, 2019; Gilotte et al., 2018; Gruson et al., 2019; Jeunen et al., 2024b). Additional to performance evaluation, these methods open up avenues for learning and fairness in recommendation—see Joachims et al. (2021) for a broad overview.
@@
Our work aims to close this gap. We derive a general expression for action propensities under TS, reducing to an integral that involves the probability density functions (PDFs) and cumulative distribution functions (CDFs) for the action-specific reward distributions. For the most common settings of (log)normal distributions for continuous outcomes and Beta distributions for binary outcomes, we show that this integral reduces to an efficiently computable multivariate normal CDF or an analytically computable expression, respectively. We show that, under a common linearity assumption on the last layer of the reward model (Chapelle and Li, 2011; Watson et al., 2021), a normal distribution on model parameters directly translates to a distribution over model outcomes (i.e. rewards) (Zhang et al., 2021), enabling efficient propensity calculations. These novel derivations open up a range of potential use-cases where either or both the logging and target policies make use of TS or related action-selection procedures.
We empirically validate our theoretical contributions by showcasing the efficacy of offline evaluation methods that would not have been feasible before, through reproducible experiments leveraging the Open Bandit Pipeline (OBP) (Saito et al., 2021).
@@ 2.1. Counterfactual or Off-Policy Estimation
Whilst this estimator is unbiased, its variance is often problematic, and a large literature on off-policy estimation has focused on variance reduction. Some trade in variance for a small bias (Ionides, 2008; Swaminathan and Joachims, 2015), others leverage additive control variates (Gupta et al., 2024b; Dudík et al., 2014; Vlassis et al., 2019; Farajtabar et al., 2018; Su et al., 2020; Jeunen and Ustimenko, 2024).
Several related works have leveraged variants of these estimators to achieve empirical success in recommender systems evaluation (Yang et al., 2018; Gilotte et al., 2018; Gruson et al., 2019; Jeunen et al., 2024b; Jeunen, 2023), as well as learning (Jeunen et al., 2024a; Jeunen and Goethals, 2021a; Jeunen et al., 2020; Chen et al., 2019; Ma et al., 2020).
@@ 3. Methodology & Contributions
Intuitively, for every possible value of $r_{1}$, the reward estimates for all other actions need to be lower for $a_{1}$ to be selected. In the general case, this integral does not have a simple analytical solution. In what follows, we derive efficiently computable or closed-form solutions for common families of reward distributions: (log)normal distributions for continuous outcomes, and Beta distributions for binary rewards.
@@ 3.1. Continuous Rewards: Normal Distributions
(6) and $\text{and~{}}\bm{V}=\begin{bmatrix}\sigma_{2}^{2}+\sigma_{1}^{2}&\dots&\sigma_{1}^{2}\\ \vdots&\ddots&\vdots\\ \sigma_{1}^{2}&\dots&\sigma_{n}^{2}+\sigma_{1}^{2}\\ \end{bmatrix}.$
Note that this multivariate Gaussian CDF is an integral without a known analytical expression in and of itself—but one that is widespread, and where many accurate and efficient implementations are available to evaluate it (Virtanen et al., 2020). Similar quantities appear in the literature on PAC-Bayesian policy learning (Sakhi et al., 2023).
@@ 3.1.1. Uncertainty on Model Parameters
Traditional TS places uncertainty on model parameters instead of model output. That is, instead of directly sampling outputs as $\hat{r}_{a}\sim f_{\theta}(a,x)$, we sample model parameters $\tilde{\theta}\sim\mathsf{P}(\Theta)$ and simply compute reward estimates as $\hat{r}_{a}=f_{\tilde{\theta}}(a,x)$, before picking the maximal one. It should be noted that in many common practical instantiations, these are equivalent. If we assume the parameters over which to place uncertainty $\theta\in\mathbb{R}^{d}$ to correspond to weights in a linear model and $f(a,x)$ is a fixed model transformation (e.g. a neural network, or the identity function in the case of a linear model), we have:
@@
3.1.2. Log-Normal Distributions
Another plausible modelling choice in recommendation applications to represent continuous rewards, is to leverage log-normal distributions. One can imagine non-negative, continuous and typically skewed outcomes like gross merchandise value or dwell time; but log-normal distributions also occur more widely in reinforcement learning value distributions (Zhang et al., 2017). When $\bar{R}$ follows a $\mathcal{N}(\mu;\sigma^{2})$ distribution, then $R\coloneq e^{\bar{R}}$ follows a $\mathrm{Lognormal}(\mu;\sigma^{2})$ distribution. Because the exponential transformation is monotonic, action propensities can be computed on $\bar{R}$ itself and we can directly apply the derivation for normal distributions here as well.
@@ 4. Empirical Validation & Discussion
To aid in the reproducibility of our work, all source code to reproduce these results as well as general implementations of our methods are made available under an open-source software license at github.com/olivierjeunen/ope-ts-recsys-2025.
@@ Acknowledgements.
The author would like to thank Duane Rich for valuable comments on earlier versions of this manuscript, and Marcelo Hartmann for insights regarding the case of (log)normal distributions.
@@ Appendix B Experimental Setup
All source code necessary to reproduce these results is available at github.com/olivierjeunen/ope-ts-recsys-2025.